Impedance: What is the impedance at 1500 Hz if a 100-Ω resistor, 20-mH coil, and 1.0-μF capacitor are connected in series?
Answer #1

Given

   frequency f = 1500 Hz

   resistance R = 100 Ω

   inductance L = 20 mH = 20*10-3 H

   capacitance C = 1.0 μF =1.0*10-6    

   Impedance Z = √ [R2 +( XL-XC)2]

    where XL = 2πf L, = 188.49

XC = 1/(2πfC) = 106.10

  Impedance Z = √ [100^2+(188.49-106.10)^2]

Impedance Z = 56.67 ohm